[Hallc_running] [New Logentry] Hall C RC update: Thursday, July 4

gaskelld at jlab.org gaskelld at jlab.org
Thu Jul 4 13:50:07 EDT 2019

Logentry Text:
Accelerator recovery proceeds, but estimated time of beam for physics has shifted a little.   The North Linac cryomodules should be full by ~ end of day shift today, and the South Linac tomorrow.  Beam recovery starts immediately after. On the White Board schedule, "restore" should last through Sunday OWL, with tune beam in the halls during day shift Sunday.

Despite the shift in expected beam start, we will still plan on condensing the target Friday in order to be ready for beam on Sunday.

Aside from target operators, shifts for Friday and Saturday OWL and DAY are cancelled. Saturday SWING and Sunday OWL are on Standby.  Target operators should start reporting Friday, DAY shift.

There will be no RC meeting today. We will meet tomorrow at the usual time.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3698663
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190704/0dad7e90/attachment.html>

More information about the Hallc_running mailing list